# CURE: Controlled Unlearning for Robust Embeddings – Mitigating Conceptual Shortcuts in Pre-Trained Language Models

### v1

@@ Abstract
CURE: Controlled Unlearning for Robust Embeddings – Mitigating Conceptual Shortcuts in Pre-Trained Language Models
@@ 1 Introduction
With the rapid advancement of artificial intelligence, pre-trained language models (PLMs) have been widely adopted across various domains, including education, healthcare and e-commerce (Devlin et al., 2019; Radford et al., 2019; Touvron et al., 2023a). A predominant strategy for applying these models is fine-tuning, where a PLM is further adapted to task-specific data, aiming to enhance its performance or better align with human intent (Ouyang et al., 2022). However, fine-tuning often exposes models to dataset biases, leading to shortcuts—spurious correlations between features and labels (He et al., 2019). For instance, Zhou et al. (2024) demonstrated that on the Yelp dataset (Zhang et al., 2015a), a LLaMA2-based (Touvron et al., 2023b) sentiment classifier mistakenly associated the concept of “food” with a “positive” label. These fragile dependencies not only limit the robustness of PLMs but also pose significant risks. In medical diagnosis, a biased detector might incorrectly associate certain biological attributes with diseases, leading to inaccurate predictions (Jiménez-Sánchez et al., 2023). Similarly, in automated recruitment systems, a shortcut may result in a favor to applicants with certain demographic attributes, exacerbating fairness problem. In Figure 1, we present an example where the classifier incorrectly associates the concept of food with positive sentiment.
Contemporary debiasing research primarily focus on two strategies: (1) modifying shortcut-inducing terms in training data (Wen et al., 2022; Yang et al., 2024b), and (2) generating counterfactual samples (Chen et al., 2023; Zhou et al., 2024) via large language models (LLMs). However, both approaches suffer from notable limitations. Lexical modification requires prior knowledge of shortcut-inducing terms, which is often challenging to obtain (Kaushik and Lipton, 2018). Moreover, its effectiveness is restricted to lexical shortcuts rather than conceptual biases. On the other hand, LLM-based counterfactual generation is computationally expensive and increases training costs significantly. While LLM-free counterfactual generation still relies on prior knowledge (Xu et al., 2023a), making it similarly constrained.
@@ 2 Related Work
Addressing spurious correlations in PLMs has become a critical research focus, as these correlations can lead to biased and unreliable predictions, limiting model robustness and fairness. Traditional works have explored various strategies to mitigate these issues, including causal inference techniques (Wang et al., 2022), adversarial training (Sagawa* et al., 2020), and data augmentation methods designed to reduce model reliance on spurious features (Kaushik et al., 2021). Additionally, approaches leveraging counterfactual reasoning (Xu et al., 2023b) have shown promise in improving fairness and robustness in LLMs. These advancements collectively contribute to a growing body of research aimed at developing more reliable and ethically sound language models.
@@ 2.1 General Approaches to Addressing Spurious Correlations
Kumar et al. (2019) addresses the challenge of models learning spurious topical shortcuts instead of relevant features in tasks like native language identification. They introduce an adversarial model to demote these latent topical confounds using log-odds ratios, guiding the model to focus on stylistic rather than topic-based features. Yaghoobzadeh et al. (2019) enhance robustness by fine-tuning models on “forgettable” examples that models initially misclassified. Stacey et al. (2020) tackle the issue of natural language inference (NLI) models relying on superficial hypothesis patterns by using an ensemble of adversarial classifiers. Wang and Culotta (2020) propose using treatment effect estimation to distinguish genuine correlations from spurious ones, such as associating “Spielberg” with positive sentiment in movie reviews. Wang et al. (2021) extend this concept with an automated framework using interpretability techniques, cross-dataset stability, and knowledge-aware perturbation to identify spurious tokens at scale. Tu et al. (2020) explores how pre-trained models like BERT handle spurious correlations, finding that they improve robustness by generalizing from minority counterexamples. The authors propose using multi-task learning (MTL) with auxiliary tasks to enhance robustness when these counterexamples are scarce. Du et al. (2022) propose the Less-Learn-Shortcut (LLS) which down-weighs examples with high correlations between specific words and labels. Xu et al. (2023b) present a counterfactual debiasing approach that balances predictions between claim-only and claim-evidence models to reduce bias associated with claim patterns. While these studies primarily address general spurious correlations, recent research has started focusing on spurious correlations at the concept level.
@@ 3.3.2 Concept-Irrelevant Content Maintenance
The semantic features are often entangled with each other (Dai et al., 2019a). As a result, although the content extractor $f_{\phi}$ solely aims at filtering out conceptual information, it is crucial to ensure that the concept-irrelevant information remain complete. Inspired by back-translation in machine translation (Sennrich et al., 2016), we construct a reversal network $\hat{\phi}$, with the same architecture as $f_{\phi}$. $\hat{\phi}$ is designed to reconstruct $x$ from $f_{\phi}(x)$, ensuring that the mapping function $f_{\phi}$ does not excessively lose the concept-irrelevant information. We first freeze $\phi$, then use the following loss to train $\hat{\phi}$:
@@ 4.1 Experimental Setup
We used IMDB (Maas et al., 2011) and Yelp (Zhang et al., 2015b) datasets. The IMDB movie review dataset is a binary sentiment analysis dataset, which consists of 50,000 positive or negative reviews from the Internet Movie Database. The Yelp dataset is provided by the Yelp Dataset Challenge, contains business reviews labeled with ratings ranging from 0 to 4 (Zhang et al., 2015b). We used the version that was cleaned and organized by Dai et al. (2019b).

### v2

# CURE: Controlled Unlearning for Robust Embeddings — Mitigating Conceptual Shortcuts in Pre-Trained Language Models

@@ Abstract
CURE: Controlled Unlearning for Robust Embeddings — Mitigating Conceptual Shortcuts in Pre-Trained Language Models
@@ 1 Introduction
With the rapid advancement of artificial intelligence, pre-trained language models (PLMs) have been widely adopted across various domains, including education, healthcare and e-commerce (Devlin et al., 2019; Radford et al., 2019; Touvron et al., 2023a). A predominant strategy for applying these models is fine-tuning, where a PLM is further adapted to task-specific data, aiming to enhance its performance or better align with human intent (Ouyang et al., 2022). However, fine-tuning often exposes models to dataset biases, leading to shortcuts—spurious correlations between features and labels (He et al., 2019). For instance, Zhou et al. (2024) demonstrated that on the Yelp dataset (Zhang et al., 2015), a LLaMA2-based (Touvron et al., 2023b) sentiment classifier mistakenly associated the concept of “food” with a “positive” label. These fragile dependencies not only limit the robustness of PLMs but also pose significant risks. In medical diagnosis, a biased detector might incorrectly associate certain biological attributes with diseases, leading to inaccurate predictions (Jiménez-Sánchez et al., 2023). Similarly, in automated recruitment systems, a shortcut may result in a favor to applicants with certain demographic attributes, exacerbating fairness problem. In Figure 1, we present an example where the classifier incorrectly associates the concept of food with positive sentiment.
Contemporary debiasing research primarily focus on two strategies: (1) modifying shortcut-inducing terms in training data (Wen et al., 2022; Yang et al., 2024b), and (2) generating counterfactual samples (Chen et al., 2023; Zhou et al., 2024) via large language models (LLMs). However, both approaches suffer from notable limitations. Lexical modification requires prior knowledge of shortcut-inducing terms, which is often challenging to obtain (Kaushik and Lipton, 2018). Moreover, its effectiveness is restricted to lexical shortcuts rather than conceptual biases. On the other hand, LLM-based counterfactual generation is computationally expensive and increases training costs significantly. While LLM-free counterfactual generation still relies on prior knowledge (Xu et al., 2023), making it similarly constrained.
@@ 2 Related Work
Addressing spurious correlations in PLMs has become a critical research focus, as these correlations can lead to biased and unreliable predictions, limiting model robustness and fairness. Traditional works have explored various strategies to mitigate these issues, including causal inference techniques (Wang et al., 2022), adversarial training (Sagawa* et al., 2020), and data augmentation methods designed to reduce model reliance on spurious features (Kaushik et al., 2021). Additionally, approaches leveraging counterfactual reasoning (Xu et al., 2023) have shown promise in improving fairness and robustness in LLMs. These advancements collectively contribute to a growing body of research aimed at developing more reliable and ethically sound language models.
@@ 2.1 General Approaches to Addressing Spurious Correlations
Kumar et al. (2019) addresses the challenge of models learning spurious topical shortcuts instead of relevant features in tasks like native language identification. They introduce an adversarial model to demote these latent topical confounds using log-odds ratios, guiding the model to focus on stylistic rather than topic-based features. Yaghoobzadeh et al. (2019) enhance robustness by fine-tuning models on “forgettable” examples that models initially misclassified. Stacey et al. (2020) tackle the issue of natural language inference (NLI) models relying on superficial hypothesis patterns by using an ensemble of adversarial classifiers. Wang and Culotta (2020) propose using treatment effect estimation to distinguish genuine correlations from spurious ones, such as associating “Spielberg” with positive sentiment in movie reviews. Wang et al. (2021) extend this concept with an automated framework using interpretability techniques, cross-dataset stability, and knowledge-aware perturbation to identify spurious tokens at scale. Tu et al. (2020) explores how pre-trained models like BERT handle spurious correlations, finding that they improve robustness by generalizing from minority counterexamples. The authors propose using multi-task learning (MTL) with auxiliary tasks to enhance robustness when these counterexamples are scarce. Du et al. (2022) propose the Less-Learn-Shortcut (LLS) which down-weighs examples with high correlations between specific words and labels. Xu et al. (2023) present a counterfactual debiasing approach that balances predictions between claim-only and claim-evidence models to reduce bias associated with claim patterns. While these studies primarily address general spurious correlations, recent research has started focusing on spurious correlations at the concept level.
@@ 3.3.2 Concept-Irrelevant Content Maintenance
The semantic features are often entangled with each other (Dai et al., 2019). As a result, although the content extractor $f_{\phi}$ solely aims at filtering out conceptual information, it is crucial to ensure that the concept-irrelevant information remain complete. Inspired by back-translation in machine translation (Sennrich et al., 2016), we construct a reversal network $\hat{\phi}$, with the same architecture as $f_{\phi}$. $\hat{\phi}$ is designed to reconstruct $x$ from $f_{\phi}(x)$, ensuring that the mapping function $f_{\phi}$ does not excessively lose the concept-irrelevant information. We first freeze $\phi$, then use the following loss to train $\hat{\phi}$:
@@ 4.1 Experimental Setup
We used IMDB (Maas et al., 2011) and Yelp (Zhang et al., 2015) datasets. The IMDB movie review dataset is a binary sentiment analysis dataset, which consists of 50,000 positive or negative reviews from the Internet Movie Database. The Yelp dataset is provided by the Yelp Dataset Challenge, contains business reviews labeled with ratings ranging from 0 to 4 (Zhang et al., 2015). We used the version that was cleaned and organized by Dai et al. (2019).